2. The physician's order reads: "Ancef 2 gram in 100 mL D5W, IVPB, to be infused over 20 minutes." The drop factor is 15 gtts/mL. How many drips per minute will the client receive? ____________________gtts/minute (Round to whole number)
Answer #1

ANSWER

2. Drips per minute of drug recevied by the client.

         Formula

         * Drips per min (gtts / min) = mL / hr * drop factor / 60 minutes.

Here,

         * mL / hr

                 * Total number of mL / Infusion time = X mL / 60 min.

                     100 mL / 20 min = X mL / 60 min

                     100 * 60 = X * 20

                                          X = 6000 / 20

                                             = 300 ml / hr.

           * Drop factor = 15 gtt / ml.

So,

         * Drips per min ( gtt / min ) = 300 * 15 / 60

                                                      = 4500 / 60

                              = 75 gtts / min.